Sara is 6 years older than 3 times age of her son Tom. the difference in Sarah's age and Tom's age is
28 years
What's Sarah and Tom's age

Answer by lwsshak3(11628)

Sara is 6 years older than 3 times age of her son Tom. the difference in Sarah's age and Tom's age is
28 years
What's Sarah and Tom's age
let x=Tom's age
3x+6=Sarah's age
2x+6-x=28
x=22
3x+6=66+6=72
Tom's age=22
3x+6=Sarah's age=72